# GraphCLIP: Enhancing Transferability in Graph Foundation Models for Text-Attributed Graphs

### v1

@@ 1. Introduction
Our proposed GraphCLIP framework. We develop GraphCLIP to address the challenges above, which learns graph foundation models with robust cross-domain zero-shot transferability from a novel self-supervised contrastive graph-summary pretraining technique. Specifically, we leverage the open LLM, QWen2-72B (Yang et al., 2024a), to generate graph-related summary texts of subgraphs, capitalizing on their powerful summarization abilities. Utilizing generated large-scale graph-summary pairs, we train a cross-domain graph foundation model through designed self-supervised contrastive graph-summary pretraining (addressing Challenge (i)). Considering the necessity of out-of-domain generalization across graphs, we introduce invariant learning during pretraining to capture invariant features, thereby enhancing out-of-domain generalization. After pretraining, GraphCLIP can be applied directly to target data without fine-tuning, demonstrating strong zero-shot capability on both in-domain and cross-domain graph data (addressing Challenge (ii)). For few-shot scenarios, we propose a novel graph prompt tuning approach aligned with our pretraining objective, reducing catastrophic forgetting (Shi et al., 2024) and minimizing learning costs (Sun et al., 2022; Liu et al., 2023b), ensuring excellent performance in few-shot settings. Extensive experiments show that GraphCLIP exhibits strong zero-shot performance across various in-domain and cross-domain target datasets. In few-shot settings, GraphCLIP with prompt tuning outperforms previous state-of-the-art methods. Additionally, the universality of our method is demonstrated through evaluation across various downstream tasks.
@@ 2.1. Notations
In this work, we focus on Text-Attributed Graphs, which incorporate raw text information for each node. Formally, given a text-attributed graph $G=\{\mathcal{V},\{T_{n}\}_{n=1}^{N},A\}$, where $\mathcal{V}$ denotes the node set with $|\mathcal{V}|=N$ instances, $T_{n}\in\mathcal{T}^{L_{n}}$ represents the raw text for node $n\in[1,2\ldots,N]$, $\mathcal{T}$ is the token dictionary, and $L_{n}$ is the sequence length, $A\in\mathbb{R}^{N\times N}$ denotes the adjacency matrix. To enhance scalability, a sampling function $\Gamma(\cdot)$ is applied to a large graph to derive a set of small ego-graphs =$\mathcal{I}=\{G_{n}\}_{n=1}^{N}$, where $G_{n}$ represents the subgraph centered on node $n\in[1,2\ldots,N]$.
@@ 3.2.2. Summary Encoding
3.2.3. Graph-Summary Contrastive Pretraining
@@
$-\underset{G,S\sim\mathbb{P}}{\mathbb{E}}\mathbb{E}_{\tau_{\alpha},\tau_{\beta}\sim\pi^{2}}\left\|g_{\theta}\left(\tau_{\alpha}\left(G\right)\right)-f_{\phi}\left(\tau_{\beta}(S)\right)\right\|^{2}$
$+\underset{S\sim\mathbb{Q}_{S}}{\mathbb{E}}\log\underset{G^{\prime}\sim\mathbb{Q}_{G}}{\mathbb{E}}\mathbb{E}_{\tau^{\prime}\sim\pi}\left[e^{\left\|f_{\phi}\left(\tau_{\beta}\left(S\right)\right)-g_{\theta}\left(\tau^{\prime}(G^{\prime})\right)\right\|^{2}}\right],$
@@ 4.2. Baselines
To evaluate the effectiveness of GraphCLIP, we compare it against 16 baselines, which include: (i) eight LLM-only methods (without modeling structural information of the graphs), i.e., BERT (Kenton and Toutanova, 2019), SBERT (Reimers and Gurevych, 2019), DeBERTa (He et al., 2020b), E5 (Wang et al., 2022), Qwen2-7B-Insturct (Yang et al., 2024a), Qwen2-72B-Insturct (Yang et al., 2024a), LLaMA3.1-8B-Instruct (Vavekanand and Sam, 2024), and LLaMA3.1-Insturct-70B (Vavekanand and Sam, 2024), (ii) 4 state-of-the-art TAG methods, i.e., GraphGPT (Tang et al., 2024), LLaGA (Chen et al., 2024b), OFA (Liu et al., 2024a), and ZeroG (Li et al., 2024b), and (iii) 4 self-supervised graph algorithms applied to TAGs, i.e., DGI (Velickovic et al., 2019), GRACE (Zhu et al., 2020), BGRL (Thakoor et al., 2021) and GraphMAE (Hou et al., 2022). Specifically, to assess the zero-shot performance of discriminative LLMs and self-supervised graph algorithms on target data, we will use the cosine similarity between node embeddings and label embeddings for predictions. For generative LLM methods, we leverage their generative capabilities to estimate their zero-shot performance. Detailed descriptions of these baselines can be found in the Appendix E.
@@ 4.3.1. Node Classification
Lastly, we observe that other self-supervised graph methods underperform across most target datasets because they do not effectively align graph and text modalities, thus failing to leverage powerful off-the-shelf text encoders.

### v2

@@ 1. Introduction
Our proposed GraphCLIP framework. We develop GraphCLIP to address the challenges above, which learns graph foundation models with robust cross-domain zero-shot transferability from a novel self-supervised contrastive graph-summary pretraining technique. Specifically, we leverage the open LLM, QWen2-72B (Yang et al., 2024a), to generate graph-related summary texts of subgraphs, capitalizing on their powerful summarization abilities. Utilizing generated large-scale graph-summary pairs, we train a cross-domain graph foundation model through designed self-supervised contrastive graph-summary pretraining (addressing Challenge (i)). Considering the necessity of out-of-domain generalization across graphs, we introduce invariant learning during pretraining to capture invariant features, thereby enhancing out-of-domain generalization. After pretraining, GraphCLIP can be applied directly to target data without fine-tuning, demonstrating strong zero-shot capability on both in-domain and cross-domain graph data. For few-shot scenarios, we propose a novel graph prompt tuning approach aligned with our pretraining objective, reducing catastrophic forgetting (Shi et al., 2024) and minimizing learning costs (Sun et al., 2022; Liu et al., 2023b), ensuring excellent performance in few-shot settings (addressing Challenge (ii)). Extensive experiments show that GraphCLIP exhibits strong zero-shot performance across various in-domain and cross-domain target datasets. In few-shot settings, GraphCLIP with our designed prompt tuning outperforms previous state-of-the-art methods. Additionally, the universality of our method is demonstrated through evaluation across various downstream tasks.
@@ 2.1. Notations
In this work, we focus on Text-Attributed Graphs, which incorporate raw text information for each node. Formally, given a text-attributed graph $G=\{\mathcal{V},\{T_{n}\}_{n=1}^{N},A\}$, where $\mathcal{V}$ denotes the node set with $|\mathcal{V}|=N$ instances, $A\in\mathbb{R}^{N\times N}$ denotes the adjacency matrix, $T_{n}\in\mathcal{T}^{L_{n}}$ represents the raw text for node $n\in[1,2\ldots,N]$, $\mathcal{T}$ is the token dictionary, and $L_{n}$ is the sequence length. Node attributes $X$ are derived by applying SBERT (Reimers and Gurevych, 2019) to the raw text information. To enhance scalability, a sampling function $\Gamma(\cdot)$ is applied to a large graph to derive a set of small ego-graphs =$\mathcal{I}=\{G_{n}\}_{n=1}^{N}$, where $G_{n}$ represents the subgraph centered on node $n\in[1,2\ldots,N]$.
@@ 3.2.2. Summary Encoding
3.2.3. Contrastive Graph-Summary Pretraining
@@
$\underset{G,S\sim\mathbb{P}}{\mathbb{E}}\mathbb{E}_{\tau_{\alpha},\tau_{\beta}\sim\pi^{2}}\left\|g_{\theta}\left(\tau_{\alpha}\left(G\right)\right)-f_{\phi}\left(\tau_{\beta}(S)\right)\right\|^{2}$
$-\underset{S\sim\mathbb{Q}_{S}}{\mathbb{E}}\log\underset{G^{\prime}\sim\mathbb{Q}_{G}}{\mathbb{E}}\mathbb{E}_{\tau^{\prime}\sim\pi}\left[e^{\left\|f_{\phi}\left(\tau_{\beta}\left(S\right)\right)-g_{\theta}\left(\tau^{\prime}(G^{\prime})\right)\right\|^{2}}\right],$
@@ 4.2. Baselines
To evaluate the effectiveness of GraphCLIP, we compare it against 17 baselines, which include: (i) eight LLM-only methods (without modeling structural information of the graphs), i.e., BERT (Kenton and Toutanova, 2019), SBERT (Reimers and Gurevych, 2019), DeBERTa (He et al., 2020b), E5 (Wang et al., 2022), Qwen2-7B-Insturct (Yang et al., 2024a), Qwen2-72B-Insturct (Yang et al., 2024a), LLaMA3.1-8B-Instruct (Vavekanand and Sam, 2024), and LLaMA3.1-Insturct-70B (Vavekanand and Sam, 2024), (ii) 4 state-of-the-art TAG methods, i.e., GraphGPT (Tang et al., 2024), LLaGA (Chen et al., 2024b), OFA (Liu et al., 2024a), and ZeroG (Li et al., 2024b), and (iii) 5 self-supervised graph algorithms applied to TAGs, i.e., DGI (Velickovic et al., 2019), GRACE (Zhu et al., 2020), BGRL (Thakoor et al., 2021), GraphMAE (Hou et al., 2022) and G2P2 (Wen and Fang, 2023). Specifically, to assess the zero-shot performance of discriminative LLMs and self-supervised graph algorithms on target data, we will use the cosine similarity between node embeddings and label embeddings for predictions. For generative LLM methods, we leverage their generative capabilities to estimate their zero-shot performance. Details of these baselines can be found in the Appendix E.
@@ 4.3.1. Node Classification
Lastly, we observe that four self-supervised graph methods, i.e., DGI, GRACE, BGRL, and GraphMAE, consistently underperform across most target datasets due to their lack of effective alignment between graph and text modalities, thereby failing to leverage powerful off-the-shelf text encoders. Furthermore, G2P2 uses a text encoder as an aligner; however, it relies on raw text as input for the text modality, which lacks comprehensive graph-level descriptions, resulting in limited transferability.
@@ Appendix E Baselines
G2P2 (Wen and Fang, 2023) proposes graph-grounded pre-training and prompting to boost low-resource text classification.